[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Feb 10, 2016

Christopher Mello - NOAA Federal christopher.mello at noaa.gov
Wed Feb 10 19:53:40 UTC 2016


Update...
Chris Mello noticed a problem during some loading some METARS on the NH1
and NH2 domains in D2D while doing some developmental work .
ticket number 722250. TSB is working with NCF to solve the problem.


Chris

On Wed, Feb 10, 2016 at 7:40 AM, Mark DeMaria - NOAA Federal <
mark.demaria at noaa.gov> wrote:

> Here is a summary of today's computer operations.
>
> Mark
>
> -----
>
> TAFB:
>
> TSB is continuing to work on getting ftp gate access from the debout
> standing desk N-AWIPS system. The required info on the system was provided
> to NCO yesterday to open a ticket with OPS-Net to get this done.
>
> NCO is finished configuring the hardware replacement of the derecho
> N-AWIPS system at the other standing desk. The new derecho hardware will be
> swapped at the standing desk at a time that is convenient for TAFB.
>
> HSU:
>
> No new issues.
>
> Other:
>
> The relocation of the SBN AWIPS cable yesterday at about 2:30 PM for the
> roof repair went well. The GOES-West antenna cable was moved at the same
> time. There was only minor loss of data during the outage, which lasted
> about 1/2 hour.
>
> There was a short internet outage around noon yesterday. TSB will follow
> up with NCO on the cause of that.
>



-- 
Chris Mello
National Hurricane Center
TSB
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160210/41bc4c03/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list